Determine the point estimate of the population​ proportion, the margin of error for the following confidence​ interval, and the number of individuals in the sample with the specified​ characteristic, x, for the sample size provided. lower bound = 1.46, upper bound = 0.634 n= 1000
Answer #1

point estimate = (lower bound+ upper bound)/2

here lower bound = 0.146 {1.46 > 1 , hence you mistyped}

upper bound = 0.634

p^= ( 0.146 + 0.634)/2 = 0.39

margin of error = (upper bound - lower bound)/2

=(0.634 - 0.146)/2

= 0.244